# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

### v2

@@ II.3 Accumulated Leakage Test
Once the 2DEG is accumulated, the accumulated leakage test is initiated. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
II.4 Screening and Reservoir Characterization
@@
The global accumulation voltage does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage in Fig. 1(b).
@@
With the 2DEG globally accumulated using the voltages selected from Sec. II.2, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig. 2(c). BATIS uses these pinch-off curves to determine the operational configuration for reservoir gates that ensures proper 2DEG accumulation and establishes bounds for operating the screening gates.
@@
The signals for screening and reservoir gates are preprocessed as described in Sec. II.2. Once cleaned, the derivative of the current vs. voltage signal, , is fit to the derivative of the phenomenological function given in Eq. (1). The pinch-off, guaranteed isolation, and operating voltages are defined following Eq. (2), with , , and , respectively. The operating voltage for the central screening gate is handled separately as it has to satisfy two constraints: ensuring that all current channels can be formed without merging. By default, . If at the subsequent state, a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
@@
The typical failure edge cases for this stage include too much depletion and too much accumulation; see the second row, the right column of Fig. 2(c). If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. On the other hand, if the channel is too accumulated, it is not possible to pinch it off. However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination. Thus, the no change in current at this stage has to be caused either by an issue with the instrument or a broken gate. To check whether this is the case, the algorithm executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
II.5 Forming 1D Current Channels
@@
The 1D current channels are formed by selectively removing 2DEG from certain regions of the device. It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along the desirable path, as seen in the top row of Fig. 2(d).
@@
Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots) [31]. If the finger gates are not accumulated enough, this region will not form. Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset. Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate). BATIS uses the channel isolation (VI) and the operating points (VO) to prepare the device for characterizing the remaining screening gates by sweeping from VI past VP. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
The plot in the left panel in the second row Fig. 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left. BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions. The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to Ref. [32]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point V for , among numerous candidates in the operating region.
If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig. 2(d), BATIS illuminates the device with the finger gates voltages set to be more negative than the most recent illumination. This adjustment can be found in the configuration file, we used 400 . After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
II.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs (VP) can be measured. Ideally, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend consistent with Eq. 1, eventually cutting off the current in the respective channels following Eq. (2) with . An idealized representation of such a curve is visualized in the left panel in the second row in Fig. 2(e).
The bottom panel in the third row of Fig. 2(e) shows a typical curve for these measurements. As the voltage is reduced, the curve begins to exhibit the characteristic Coulomb blockade oscillations typical of QD devices [13]. In principle, such oscillations could complicate any fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage (the reverse of how the signal was measured) to evade the small deviations of background noise, whilst capturing the first significant turn-on.
III Experimental Demonstration
@@
We now demonstrate BATIS on a 1D quadruple Si/SixGe1-x QD device, nominally identical to the one shown in Fig. 1(a) with 24 gates. The first layer of gates is made up of four screening gates, Si for , shown in Fig. 1(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in Fig. 1(a) in blue, while the plungers are shown in red. The final layer consists of nine barrier gates, Bi for , shown in Fig. 1(a) in gray.
@@
The goal of BATIS is to form three current channels in the quad-QD device as it learned from the gate layout in the configuration, as indicated with white arrows in Fig. 3(a). BATIS brings a grounded but not characterized device from Fig. 1(a) to a fully operational device in Fig. 1(d). In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterizations.
@@
First Leakage Test. BATIS begins by performing a device leakage test. For the quad-QD device used in our demonstration, the sample mount holding it contains connections, of which are used for instrument channels to control individual gates and ohmics and an additional wired channels are unused for this device. Thus, the size of the leakage matrix is to examine all possible leakage states. BATIS first checks the diagonal elements of the leakage matrix. If the resistance of every connection exceeds 25, BATIS declares that there is no leakage in the device.
If any of the diagonal elements falls below the 25 threshold, the device is declared leaking. Depending on where the current leak occurs, it is sometimes possible to recover a leaking device. Thus, rather than terminating, BATIS continues to fill in the leakage matrix for diagnostic purposes. Due to the symmetry of the measurements, the leakage matrix is also symmetric and only the lower triangular part of the matrix needs to be completed. Moreover, passing the threshold test by a diagonal component automatically guarantees that the entire column corresponding to that component also passes the test. Thus, only the columns corresponding to the elements below 25 need to be determined. BATIS continues to fill the leakage matrix one column at a time. In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
@@
The leakage test needs to be executed only once per cooldown if under fully automatic control. This is because BATIS has built-in safety checks during voltage choice selection to avoid exceeding safe voltage breakdowns between neighboring gates for all measurements.
| Voltages | Run 1 | Run 2 | Run 3 | Run 4 | |
| I1 | VP () | ||||
| Vmax () | |||||
| I2 | VP () | ||||
| Vmax () | |||||
| I3 | VP () | ||||
| Vmax () |
@@
Global Accumulation. Next, BATIS performs the global accumulation stage. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring between and . By design, all gates are grounded during the first two illuminations. The first illumination is intended to completely reset the device while the second is to verify that all instruments work as intended. Because BATIS still did not find the gates in the desired range, it chose to apply voltages to all screening and reservoir gates for the third illumination. The voltages required to achieve turn-ons at around , calculated following the protocol in Ref. [21], were set to for gates associated with channel I1 and I2, and for gates associated with channel I3. The central screening gate, S2, was left grounded during the illumination.
@@
The fourth iteration of global turn-on measurement was successful. The summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in the left panel in Fig. 1(c), where the pinch-offs on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
@@
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is measured, following the same procedure as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since the ohmic connections are excluded, the resulting matrix has five fewer indexes. As none of the connections leaked, only measurements following the matrix diagonal were taken.
@@
Screening and Reservoir Characterization. The next stage in BATIS involves the characterization of the screening and reservoir gates. The results of these characterizations for all tested gates are presented in Fig. 3(b). Unlike the other screening and reservoir gates, S2 needs to be measured simultaneously on channels I1, I2, and I3 since it borders all of the current channels. Because of this, the data for the S2 pinch-off can be found in Fig. 3(c) as it needs to be organized by channel.
@@
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature from the illumination from earlier [21]. We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long as compared to all other screening gates. This is likely due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to be more confident in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 3(b) is truncated to be consistent with the remaining measurements for visualization purposes.
The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2. A device in this configuration prepared for the channel accumulation stage is shown in Fig. 1(c).
Forming 1D Current Channels. With all screening and reservoir gates characterized, BATIS proceeds to form the 1D current channels. The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig. 3(a). BATIS identified three different operating points, one for each 1D current channel calculated from the simultaneous channel accumulation in Fig. 3(c). The summary of the screening and finger gates configuration for each operating point is presented in Table 3, and a collapsed zero cross-talk visualization is pictured in the right panel in Fig. 1(c). The three points are consolidated into one where the voltage color for each gate is the VO indexed by the channel the gate belongs to.
As we discuss in Sec.II.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots [31, 32] are orthogonal to the measured planes shown in Fig. 3(c) and would have been measurements with the S2 axis instead. In other words, the triangle plots and the plots utilized in BATIS contain similar information on different plane cuts through the same space.
@@
The accumulation plots in Fig. 3(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in . This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents. Such a dip in the charge sensor readout is often called a shadow. Shadows increase the difficulty of detecting and forming isolated current channels. The accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some cross-talk. This is due to separation with the S2 gate which divides the 2DEG between the upper readout channels and the lower dot channel.
@@
Finger Gate Characterization. The last stage of BATIS involves finger gate characterization. This characterization is done individually for all of the finger gates on the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others. This can be seen in the raw data pictured in Fig. 3(d). The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 1(d).
IV Summary and outlook
In this paper, we propose and demonstrate BATIS, a fully autonomous procedure for bootstrapping, automated testing, and characterization of accumulation mode undoped QD devices. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure, the device geometry, the ohmic biases, and the device limits. BATIS navigates through complex high-dimensional gate space to find the final gate configuration in which all desirable current channels are formed following simple physical principles.
The goal of BATIS is to test the device’s functionality to confirm that all gates can properly pinch-off and that all current channels can be formed. Testing and characterizing an unknown QD device is essential for identifying global device states that are suitable for observing the Coulomb blockade. Importantly, the single-button-press testing and initialization process is demonstrated at , eliminating the need for a full cooldown just to learn that a device is not functional. This result is particularly important as we envision a hot large-scale spin-based quantum processor. These systems will require qubit initialization, relying on a complete and accurate device characterization. One could imagine using BATIS to probe the attributes of many devices spread across a wafer or provide device diagnostics on a portion of a large qubit processor without going to temperatures.
The flexible structure of BATIS allows for scalable characterization and initialization of any accumulation-mode linear array of QDs. It is also easily adaptable to other QD platforms. In particular, BATIS could straightforwardly mimic the behavior of algorithms proposed for nanowires and GaAs/AlxGa1-xAs by simply invoking skip logic on the steps not necessary for these devices. On the other hand, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quad-QD device considered in this work.

### v3

@@ III.3 Accumulated Leakage Test
With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
III.4 Screening and Reservoir Characterization
@@
The global accumulation voltage does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage; see Fig. 2.
@@
With the 2DEG globally accumulated, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig. 3(c). BATIS uses these pinch-off curves to determine the operating points for reservoir gates, ensuring proper 2DEG accumulation, and to determine the operating bounds for the screening gates.
@@
The derivative of the response signals for screening and reservoir gate pinch-offs are preprocessed in a similar fashion as described in Sec. III.2. The pinch-off, guaranteed isolation, and operating voltages are defined following Eq. (2), with , , and , respectively. The pinch-off occurs at roughly the inflection of the signal VP, the guaranteed isolation, VI, happens left of the pinch-off, and the operating voltages, VO, occur to the right of the pinch-off, typically on the slope. The operating voltage for the central screening gate is handled separately as it has to satisfy a secondary constraint: all current channels must be formed without merging. By default, . If at the subsequent stage a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
@@
Typical unsatisfactory operational modes for this stage include too much depletion or accumulation; see the second row, the right column of Fig. 3(c). If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. On the other hand, if the channel is too accumulated, it is not possible to pinch it off. However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination. Thus, at this stage, the constant current signal is caused by either an instrument failure or a broken gate. To verify whether this is the case, BATIS executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
III.5 Current Channel Accumulation
@@
The 1D current channels are formed by selectively modifying the 2DEG in specific regions of the device. It involves depleting charges underneath the screening gates and accumulating them underneath the plunger and barrier gates (often collectively called finger gates) such that the current flows between the screening gates along the desired path, as seen in the top row of Fig. 3(d).
@@
Determining voltages where 1D channels are properly formed is typically done iteratively, by sweeping two screening gates against each other in a 2D measurement while holding the corresponding finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots) [45]. If the finger gates are not sufficiently accumulated, this region will fail to form. Multiple 2D scans with incremented finger gate voltages for each pair of screening gates—a three-dimensional dataset—are often necessary to observe the triangle region across all required channels.
BATIS employs an alternative approach for initializing multiple 1D channels simultaneously. Rather than sweeping screening gates against each other, BATIS performs a 2D measurement of a screening gate against all the finger gates located between that gate and the opposing screening gate used to define the 1D channel (typically the central screening gate). Thus, this method efficiently captures the behavior of all channels in a single 2D dataset.
A simulation of current as a function of gate voltages is shown in Fig. 4(a), and in Fig. 4(b) a corresponding example of experimental data is shown. Notably, there are four unique regions in the landscape in Fig. 4(a): a no-current region in the bottom left, a finger-gate-dominated region in the top left (indicated with I), a screening-gate-dominated region in the bottom right (indicated with III), and a region of combined effect from both the fingers and screening gates (indicated with II). These regions all correspond to different operating modes of the device, with region I showing the proper accumulation of a 1D channel that is the required precursor to proper bootstrapping of the device; region II, in contrast, corresponds to an improper 2D channel formed under the screening and finger gates; and region III corresponds to an undesired 2D channel under the screening gate. All three of these regions are indicated in the simulation results shown in Fig. 4(c). If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig. 3(d), BATIS initiates illumination with finger gates voltages set more negative than the most recent illumination.
When the finger-gate-dominated region forms properly, BATIS automatically selects an operating point in region I, near the boundary between the current and the no-current regions. Fig. 4(d) shows a Schrodinger-Poisson simulated example with properly set operating points for all current channels. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of screening/finger gates, which makes it much more efficient than the standard approaches.
III.6 Finger Gate Characterization
Once isolated current channels are formed, finger gate pinch-offs, VP, can be measured. Naively, one might hope that, as the voltage on a given finger gate is lowered, the pinch-off curves should exhibit a smooth, sigmoidal trend like that shown in the middle row of Fig. 3(e). However, because finger gates are so small, trapped charge, e.g., in the oxide layer, often causes Coulomb blockade effects [26] to be visible, so that real devices display behavior like that shown in the third row of Fig. 3(e). In this case, as the gate voltage is made less positive, the curve exhibits characteristic Coulomb blockade oscillations typical of QD devices. To be robust against this non-ideality, BATIS uses thresholding and slope detection during analysis to determine the pinch-off voltage on a point-by-point basis, going in the direction of increasing voltage from the end of the sweep (the reverse of how the signal was measured) to avoid small deviations of background noise, while capturing the first significant turn-on behavior.
IV BATIS: Experimental Demonstration
@@
We demonstrate BATIS at 1.3 on a Si/SixGe1-x quadruple QD device with 24 gates, nominally identical to the one shown in Fig. 1(a), which has two charge sensing channels in parallel with a channel to host the four qubit QDs. The first gate layer contains four screening gates, Si for , shown in Fig. 1(a) in yellow. The next layer consists of five reservoirs, Ri for , and six plunger gates, Pi for . The reservoirs are shown in blue in Fig. 1(a), while the gates highlighted in red are the plunger gates. The final layer consists of nine barrier gates, Bi for , shown in Fig. 1(a) in gray.
@@
The goal of BATIS is to form three current channels and determine proper pinch-off gate voltages based on automated measurements and the gate layout information BATIS learns from the configuration file. This proper operation is indicated with black arrows in Fig. 1(b), bringing a grounded but not-characterized device to a fully operational mode. In what follows, we explain how BATIS navigates the execution of the quad-QD device bootstrapping, testing, and characterization.
@@
First Leakage Test. The sample mount holding the quad-QD device used here contains connections, of which are used for instrument channels to control individual gates and ohmics, and an additional wired channels are unused for this device. Thus, the size of the leakage matrix is to capture all possible leakage states. The leakage matrix is symmetric, and thus only the lower triangular part needs to be completed.
BATIS first checks the diagonal elements of the leakage matrix. If the resistance of any connection drops below 25 , BATIS declares that the device leaks. Moreover, if the threshold test fails on a diagonal component, BATIS automatically fills out the rest of the column for diagnostic purposes. In the worst case, a total of resistance measurements are necessary to complete the leakage matrix.
@@
Under fully automatic control, the leakage test needs to be executed only once per cooldown. Damage following this test is unlikely, as BATIS incorporates built-in safety checks throughout its automated voltage selection process for all measurements to prevent dielectric breakdown between neighboring gates.
| Voltages () | Run 1 | Run 2 | Run 3 | Run 4 | |
| I1 | VP | ||||
| Vmax | |||||
| I2 | VP | ||||
| Vmax | |||||
| I3 | VP | ||||
| Vmax |
@@
Global Accumulation. Next, BATIS seeks to determine gate voltages that establish global accumulation. This step generates current flow across the device without regard to which channel or channels are connected to each other. In our demonstration, the algorithm executed three rounds of illumination before the observed global turn-on values were determined to be satisfactory, occurring in the desired range of to . By design, all gates are grounded during the first two illuminations. The first illumination is intended to completely reset the device, while the second is to verify that all instruments work as intended. Because BATIS still did not find the gates in the desired range for the third illumination, voltages were applied to all screening and reservoir gates. The voltages required to achieve turn-ons in the required range are determined by BATIS following the protocol in Ref. [34]. In this case, during illumination, BATIS sets gates associated with channels I1 and I2 to , and gates associated with channel I3 to . The central screening gate, S2, was left grounded (i.e., set to 0 ) during illumination.
@@
Following the gate-biased illumination, the fourth iteration of the global turn-on measurement was successful. The summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in Fig. 5(a), where the pinch-off voltages on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
@@
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is measured, following the same procedure as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 . Since none of the connections leaked, only measurements following the matrix diagonal were taken.
@@
Screening and Reservoir Characterization The next step BATIS follows involves characterization of the screening and reservoir gates, which is important for dividing the induced 2DEG into three current channels. The results of these characterizations for all tested gates are presented in Fig. 6(b). The S2 gate borders all of the current channels and thus is not included here as it needs to be measured simultaneously on channels I1, I2, and I3. The data for the S2 pinch-off can be found in Fig. 6(c).
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature that arises in part from the gate-biased illumination described above [34]. We note that for gate S1, BATIS acquired the signal over a range, i.e., twice as long compared to all other screening gates. This is due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to ensure sufficient confidence in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 6(b) is truncated to be consistent with the remaining measurements for visualization purposes.
The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2. A device in this configuration prepared for the channel accumulation stage is shown in Fig. 5(b).
@@
Current Channel Accumulation With all screening and reservoir gates characterized and candidate operating voltages determined, BATIS proceeds to form the 1D current channels. As discussed above in Sec. III.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots are along a different set of axes (e.g., S2 and S1, instead of all finger gates and S1) than those shown in Fig. 6(c). The advantage of measurements implemented in BATIS is faster convergence to the correct operating point for three well-formed and independent current channels.
The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, as explained in Ref. [46]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final operating point V for , among the numerous candidates in the operating region.
If BATIS was unable to select an operating point after such a measurement, the next action would be to adjust all of the gates by using a 400 illumination. This adjustment can be found in the configuration file. After such an illumination, the entire device response can change because of crosstalk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig. 6(a). In this demonstration, BATIS identified three different operating points, one for each 1D current channel, based on the data shown in Fig. 6(c). The summary of the screening and finger gates configuration for each operating point can be found in Table 3, and a collapsed visualization of these gate voltages is shown in Fig. 5(c).
@@
The accumulation plots in Fig. 6(c) reveal a dip in the current exiting when the current starts to flow in , shown by the dark green shape imposed on top of the signal in . This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents. Such a dip in the charge sensor readout is often called a shadow. Shadows increase the difficulty of detecting and forming isolated current channels, yet BATIS can handle this situation. In contrast, the accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some crosstalk arising from gate-to-2DEG cross capacitances.
Finger Gate Characterization The last stage of BATIS involves the determination of pinch-off voltages for the finger gates. One of the problems with analyzing pinch-off data is the presence of spurious Coulomb blockade oscillations, as discussed above and shown in the bottom row in Fig. 3(e). Such oscillations complicate fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage, searching in the increasing voltage direction (the reverse of how the signal was measured) to capture the first significant turn-on above the noise.
@@
This pinch-off voltage determination is performed individually for all of the finger gates of the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others, as is visible in the data shown in Fig. 6(d), and the added complexity of the curves informed the algorithm described above and implemented here. The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 5(d). This voltage configuration describes the proper operating point in high-dimensional voltage space. BATIS has taken the device from an unknown initial state to a set of properly defined 1D channels with confirmed finger gate operation and a full compilation of pinch-off voltages. At this point, autotuners designed for 1D devices can be used on each individual channel, if desired [9].
V Summary and outlook
In this paper, we demonstrated an autonomous system for bootstrapping, automated testing, and characterization of gate-defined, accumulation-mode, undoped QD devices. This method, which we call BATIS, leverages knowledge from experiment and physical intuition—grounded in fundamental device physics—to electronically control navigation of the quantum device landscape. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure type, the device geometry, the ohmic biases, and the device safety limits. The method proceeds in several steps, checking for device functionality as it moves forward. If irrecoverable device flaws are discovered, BATIS reports those issues and halts. Importantly, BATIS can recover from other non-idealities, such as undesirable amounts of trapped charge arising from fabrication or accident, by using illumination under gate-bias to modify in-situ the trapped charge and thus the device performance. This capability is important for resolving issues like charge disorder or incomplete turn-on, which may prevent functional behavior in otherwise operational devices.
While much of the automated approach relies on existing experimental procedures, an innovation in BATIS is its method for 1D channel formation: rather than sweeping pairs of screening gates, BATIS performs a single 2D voltage scan between one screening gate and all intermediate finger gates. This strategy reduces the required measurement overhead to a single scan, resulting in substantial efficiency gains compared to conventional approaches.
Importantly, the single-button-press testing and characterization process is demonstrated at , eliminating the need for a full cooldown for the device initialization. This result is particularly important as work advances towards “hot”, large-scale spin-based quantum processors. BATIS can also serve as a screening tool before cooling a dilution refrigerator to reach temperatures, if cold operation is desired.